Find two numbers whose sum is 20 and whose product is 96.
Answer by checkley77(12569)

X+Y=20 OR X=20-Y
XY=96
(20-Y)Y=96
20Y-Y^2=96
Y^2-20Y+96=0
(Y-8)(Y-12)=0
Y-8=0
Y=8 ANSWER.
Y-12=0
Y=12 ANSWER.